OpenStudy (anonymous):

Can some one please help me with this question 1/m=m-34/2m^2 ?

OpenStudy (anonymous):

\[\frac{ 1 }{ m } = \frac{ m-4 }{ 2m^2 }\] You mean this? So.. what will we do? Find the value of m?

OpenStudy (anonymous):

can you explain to me how to do it like by steps ?

OpenStudy (anonymous):

yes that is what i meant find the value of m

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

cross multiply. 1(2m^2) = m(m-34) 2m^2 * 1 = m * (m - 34) 2m^2 = m^2 - 34m m^2 + 34m = 0 m(m + 34) = 0 m = 0 or m + 34 = 0 m = 0 or m = -34 since its a denom., so m = -34.

OpenStudy (anonymous):

thankyou so much

OpenStudy (anonymous):

No problem. Thanks for the medal :)